If f(x)=2x-4 find f(q+1)

Found 2 solutions by josgarithmetic, solver91311:
Answer by josgarithmetic(39618)
Answer by solver91311(24713)



if f(x) = 2x - 4, then f(1) = 2(1) - 4, f(756) = 2(756) - 4, f(theartistformerlyknownasPrince) = 2(theartistformerlyknownasPrince) - 4, and f(anything you want to put in here) = 2(anything you want to put in here) - 4

So what do YOU think f(q + 1) is equal to? Remember to simplify your answer.

John

My calculator said it, I believe it, that settles it